OpenStudy (anonymous):

Multiply: x/6y * 3y^2/11

OpenStudy (goten77):

\[\frac{ x }{ 6y } \times \frac{ 3y ^{2} }{ 11 } = \frac{ 3xy ^{2} }{ 66y }\]

OpenStudy (anonymous):

I don't think that's the answer simplified, that's not one of my choices.

OpenStudy (goten77):

it can also be simplified to

OpenStudy (goten77):

\[\frac{ xy }{ 22 }\]

OpenStudy (anonymous):

and another question, would it be the x or y variable with 0 restrictions?

OpenStudy (goten77):

they both have restrictions cause if they = 0 itel become indeterminate

OpenStudy (anonymous):

because my answer choices that match are A. \[\frac{ xy }{ 22 } x \neq0\] B. \[\frac{ xy }{ 22 } y \neq 0\]

OpenStudy (phi):

look at the original fraction. What values of x or y would cause you to divide by 0? (dividing by 0 is not allowed, because it causes an indeterminate value)

OpenStudy (anonymous):

I think y. Is that right ?